OpenStudy (anonymous):

Convert the polar equation r 2 = 2tan θ to a Cartesian equation.

OpenStudy (austinl):

Is that \[r^2=2\tan(\theta)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay the answer I have is x^2+y^2=2y

OpenStudy (anonymous):

that should be right

OpenStudy (anonymous):

Thats what I thought but the thing says its wrong.

terenzreignz (terenzreignz):

rightly so. It doesn't get any better than when your polar equation is expressed only in terms of \(r^2\) and \(\tan \theta\) Because they're readily replaceable

hartnn (hartnn):

r^2 = x^2+y^2 and tan theta = y/x so, shouldn't it be x^2+y^2 = 2y/x ?

OpenStudy (anonymous):

No My school program says its x^3+xy^2=2y, but that makes no sense at all.

OpenStudy (phi):

x= r cos A y= r sin A \[ \tan A= \frac{\sin A}{\cos A}= \frac{r \sin A}{r \cos A}= \frac{y}{x} \] \[ r^2 = x^2 + y^2 \] so you have \[ r^2 = 2 \tan A \\ x^2+y^2 = \frac{2y}{x} \\ x^3 + xy^2 = 2y \]

hartnn (hartnn):

x^2+y^2 = 2y/x got this ? then just multiply x on both sides

OpenStudy (anonymous):

okay that makes sense

OpenStudy (anonymous):

Thanks All!